Write in slope-intercept form the equation of the line that passes through (-2,5) and (4,6)
Found 2 solutions by richwmiller, lwsshak3:
Answer by richwmiller(17219)

y= 1/6 x+5 1/3

Answer by lwsshak3(11628)

Write in slope-intercept form the equation of the line that passes through (-2,5) and (4,6)
***
standard form of equation for a line: y=mx+b, m=slope, b=y-intercept
slope=∆y/∆x=(6-5)/(4-(-2))=1/6
equation: y=(1/6)x+b
solve for b using coordinates of one of the given points (4,6)
6=4/6+b
b=6-4/6=18/3-2/3=16/3
equation: y=x/6+16/3